# Enhancing Temporal Consistency in Video Editing by Reconstructing Videos with 3D Gaussian Splatting

### v3

@@ Abstract
Recent advancements in zero-shot video diffusion models have shown promise for text-driven video editing, but challenges remain in achieving high temporal consistency. To address this, we introduce Video-3DGS, a 3D Gaussian Splatting (3DGS)-based video refiner designed to enhance temporal consistency in zero-shot video editors. Our approach utilizes a two-stage 3D Gaussian optimizing process tailored for editing dynamic monocular videos. In the first stage, Video-3DGS employs an improved version of COLMAP, referred to as MC-COLMAP, which processes original videos using a Masked and Clipped approach. For each video clip, MC-COLMAP generates the point clouds for dynamic foreground objects and complex backgrounds. These point clouds are utilized to initialize two sets of 3D Gaussians (Frg-3DGS and Bkg-3DGS) aiming to represent foreground and background views. Both foreground and background views are then merged with a 2D learnable parameter map to reconstruct full views. In the second stage, we leverage the reconstruction ability developed in the first stage to impose the temporal constraints on the video diffusion model. This approach ensures the temporal consistency in the edited videos while maintaining high fidelity to the editing text prompt. We further propose a recursive and ensembled refinement by revisiting the denoising step and guidance scale used in video diffusion process with Video-3DGS. To demonstrate the efficacy of Video-3DGS on both stages, we conduct extensive experiments across two related tasks: Video Reconstruction and Video Editing. Video-3DGS trained with 3k iterations significantly improves video reconstruction quality (+3 PSNR, +7 PSNR increase) and training efficiency (×$\times 1.9$, ×$\times 4.5$ times faster) over NeRF-based and 3DGS-based state-of-art methods on DAVIS dataset, respectively. Moreover, it enhances video editing by ensuring temporal consistency across 58 dynamic monocular videos.
@@ 1 Introduction
Therefore, we first design the framework of Video-3DGS to address the complexities of dynamic monocular video scenes. One of the key distinctions of the proposed method lies in its ability to excel in “wilder” monocular video scenarios (e.g., videos with larger object motion). To this end, we initially simplify the video sequences with two decomposition strategies. First, spatial decomposition, powered by an off-the-shelf open vocabulary video segmentor, mitigates background clutter 111background containing complex regions that clutter the foreground objects (e.g., trees in the background).. Second, temporal decomposition breaks down the entire video sequence into multiple shorter video clips with overlapping frames between neighboring clips. Given these decompositions, we can effectively extract 3D points of masked foreground moving objects in each video clip by introducing MC-COLMAP. Meanwhile, the cluttered background is modeled with spherical-shaped random 3D points surrounding the pre-extracted 3D points of foreground moving objects. For each video clip, Video-3DGS utilizes two sets of 3D Gaussians (Frg-3DGS and Bkg-3DGS) to represent foreground and background 3D points, respectively. We additionally implement multi-resolution hash encoding (Müller et al., 2022) with deformation networks (Yang et al., 2023) on both Frg-3DGS and Bkg-3DGS, which can boost performance and efficiency. Last but not least, to obtain the final rendered 2D outputs from both 3D Gaussians, we adopt a straightforward and effective 2D learnable parameter map to merge the two rendered images (rendered views from Frg-3DGS and Bkg-3DGS), resulting in faithful video frame representations.
@@
2 Related Work
In this part, we briefly review the literature on Text-to-Image editing (Section 2.1), and Zero-shot Text-to-Video editing (Section 2.2). More related works about neural rendering for monocular videos are introduced in Appendix A.
2.1 Text-to-Image Editing
2.2 Zero-shot Text-to-Video Editing
@@ 3.2.1 MC-COLMAP
where $\mathcal{V}^{f}$ and $class$ denote the extracted segmentation masks of foreground moving objects Wang et al. (2021); Yu et al. (2022) and the user-guided text prompt (a required input for the segmentation network to specify the target object (Yu et al., 2023a, b)), respectively.
Temporal Decomposition The segmentation masks $\mathcal{V}^{f}$ for foreground objects, which span the entire video sequence, present a processing challenge for COLMAP due to their intricate motion dynamics. To mitigate this complexity, stemming from diverse motion patterns throughout the video sequence, we partition the video sequence into multiple shorter video clips (Kim et al., 2022; Shin et al., 2024; He et al., 2023). This division ensures that objects exhibit reduced motion within each clip, thus facilitating more manageable processing for COLMAP. Formally, we address this by splitting $\mathcal{V}^{f}$ into multiple $M$ clips $\{\mathcal{V}^{f}_{j}\}_{j=1}^{M}$.
Rather than evenly dividing the video sequence into $M$ clips, each containing $k$ frames, we introduce a progressive scheme to address potential failure cases, such as when the foreground object remains static within a clip, making the $k$ frames inadequate for point cloud extraction or registration in SfM. Specifically, we begin with the first clip initialized with the first $k$ frames (i.e., $\mathcal{V}^{f}_{1}=\{\mathbf{x}_{i}\}^{k}_{i=1}$). If the COLMAP function $R$ fails to process the clip (i.e., $status_{1}\neq$ ‘Success’), we iteratively include one additional consecutive frame into the current clip until COLMAP returns a ‘Success’ $status$. Subsequently, the next clip commences with the last frame of previous clip, also initialized with $k$ frames. This process continues until the entire video sequence is processed. We term the resulting pipeline as MC-COLMAP, which applies COLMAP in a Masked and Clipped manner. Our MC-COLMAP system (denoted as $R_{MC}$) yields multiple sets of masked and clipped 3D point clouds, along with their corresponding views from $M$ clips:
@@ 3.3 Video-3DGS (2nd Stage): Plug-and-Play Refiner for Editing Videos
Single-phase Refiner The utilization of pre-trained 3D Gaussians showcases remarkable proficiency in preserving structure and ensuring temporal consistency across video frames, making them highly suitable for video editing tasks. Initially edited video frames, denoted as $\{I^{edited}_{i}\}^{N}_{i=1}$, can be sourced from any zero-shot video editors. To address inconsistencies in style and object presence within these edited videos, we leverage pre-trained 3D Gaussians. This approach involves maintaining the original structural context by fixing the positional $(x,r,s)$ and deformation parameters ($\delta$) of both Frg-3DGS and Bkg-3DGS. Simultaneously, we adjust the color value ($SH$) and opacity ($\sigma$) parameters to align with the style of the edited frames, accomplished by minimizing the reconstruction loss between rendered images $\{I^{merged}_{i}\}^{N}_{i=1}$ and edited frames $\{I^{edited}_{i}\}^{N}_{i=1}$ for updating corresponding $SH$ and $\sigma$ as follows:
@@
Altering the color value of 3D Gaussians enables corresponding areas of rendered images to maintain consistency within each clip. Moreover, refining on edited overlapping frames between neighboring clips can ensure smooth style transition across clips, ultimately enhancing the temporal consistency over whole video frames. Given that we adopt Video-3DGS following a single-phase of video editing, Video-3DGS (2nd stage) can be appropriately referred to as a single-phase refiner.
Recursive and Ensembled Refiner In the single-phase refiner, we update the spherical harmonic coefficients ($SH$) and opacity ($\sigma$) of pre-trained Video-3DGS by minimizing the reconstruction loss between rendered images and edited video frames. However, we encountered instances where some initial edited frames had already lost the structures of the original video due to prolonged inference steps (number of denoising steps) implemented by the off-the-shelf video editor. To address this issue, the number of denoising steps ($N_{d}$) is divided into multiple phases ($N_{r}$, fixed to 2 for efficiency). We perform the single-phase refinement on edited frames from the first phase (e.g., video editing after $N_{d}$/2), resulting in intermediate refined frames. These refined frames are then used as input for DDIM inversion (as utilized in TokenFlow and RAVE) or the image encoder (as employed in Text2Video-Zero) to initialize the next phase. Subsequently, the remaining denoising steps ($N_{d}$/2) are carried out for the next phase of video editing. This is followed by a final single-phase refinement using Video-3DGS on the edited frames from this subsequent phase. We refer to this process as recursive refinement. Another critical hyperparameter in the quality of edited videos is guidance parameters in the diffusion process, such as image guidance in Text2Video-Zero, text guidance in TokenFlow, and controlnet guidance in RAVE. As shown in  Figure 3, different guidance scales lead to significant changes. To mitigate sensitivity to scale changes and integrate the advantages of each scale, we conduct an ensemble update of $SH$ and $\sigma$ with multiple edited videos from three different scales during each recursive phase. For the next phase refinement, we incorporate those multiple edited videos from the previous phase into next ones. It allows each phase to benefit from the previous edits. For simplicity, we refer to Video-3DGS (2nd stage) with Recursive and Ensembled refinement as Video-3DGS (RE) and illustrate it in Figure 5.
@@ 4.2 Video Editing
Quantitative Results We assessed Video-3DGS’s editing capability on top of those three video editors in Table 2. Video-3DGS significantly enhances temporal consistency (WarpSSIM score) across various editing scenarios spanning three different datasets. Consequently, this improvement generally yields superior final video editing results ($Q_{edit}$). It proves that Video-3DGS generally provides robust temporal editing guidance to existing video editors. We also observe that our proposed recursive and ensembled refinement can further improve both temporal consistency and video editing results on all of video editors and datasets.
@@ 5 Conclusion
  • Appendix A provides more related works.
@@
  • Appendix H discusses broader impact.
@@ Appendix A Related Work
In this part, we further introduce the literature related to the core part of Video-3DGS (1st stage): neural rendering for scene representation (Section A.1), 3D Gaussian Spatting-based representation (Section A.2), dynamic scene representation (Section A.3), and spatial temporal decomposition in neural rendering (Section A.4).
@@ A.1 Neural Rendering for Scene Representation
Building upon the advancements in Neural Radiance Fields (NeRF) (Kerbl et al., 2023) for scene representation, another innovative approach that has gained attention is the utilization of 3D Gaussian distributions to model scenes. This methodology, often referred to as “Gaussian Scene Representation”, leverages the mathematical properties of Gaussian distributions to efficiently capture the volumetric density and color information of 3D scenes. Similar to the trend in NeRF, there have been several advancements in performance and efficiency. For example, some of them apply the module used in NeRF to 3D Gaussian-based view synthesis.  (Lee et al., 2024) leverages a compact MLP to adjust the covariance of 3D Gaussians, which can effectively reconstruct sharp details from blurry images and Mip-splatting (Yu et al., 2023c) address alias issue in 3DGS as similar to  (Barron et al., 2021). Besides the advancements in quality,  (Fan et al., 2023) enhances the efficiency and compactness of 3D Gaussian Splatting for neural rendering by pruning insignificant Gaussians.
@@ A.4 Spatial Temporal Decomposition
Our proposed Video-3DGS decomposes the dynamic video spatially and temporally to represent original video effectively. Similarly, ProgressiveNeRF (Meuleman et al., 2023) proposes method of decomposing video into spatial and temporal components through progressively optimized local radiance fields. It represents a significant advancement in the field of robust view synthesis and large-scale scene reconstruction. The primary distinction between our method and the ProgressiveNeRF approach lies in how we manage camera poses and frame integration across video clips. While ProgressiveNeRF dynamically updates camera poses and continues to append frames until the camera’s pose extends to the boundary of a high-resolution, uncontracted space, our method treats each clip independently. We utilize overlapping frames solely to ensure consistency of 3D Gaussians corresponding to overlapping frames across neighboring clips, with each clip maintaining its unique camera poses. However, all clips can still be rendered onto the same image plane via a differentiable rendering with each distinctive 3D Gaussian and camera pose. Additionally, our process repeats until MC-COLMAP signals a ‘Success’ status marking a clear difference in the criteria used to determine the sufficiency of frames for each clip. Furthermore, ProgressiveNeRF uses several dependencies (e.g., depth estimation and optical flow) and many losses for better temporal decomposition. However, we only need segmentation masks and simple reconstruction losses (L1 and SSIM) as in the original 3DGS paper. We also provide extensive visual comparison with ProgressiveNeRF in this link: https://video-3dgs-progressivenerf.github.io/.

### v4

@@ Abstract
Recent advancements in zero-shot video diffusion models have shown promise for text-driven video editing, but challenges remain in achieving high temporal consistency. To address this, we introduce Video-3DGS, a 3D Gaussian Splatting (3DGS)-based video refiner designed to enhance temporal consistency in zero-shot video editors. Our approach utilizes a two-stage 3D Gaussian optimizing process tailored for editing dynamic monocular videos. In the first stage, Video-3DGS employs an improved version of COLMAP, referred to as MC-COLMAP, which processes original videos using a Masked and Clipped approach. For each video clip, MC-COLMAP generates the point clouds for dynamic foreground objects and complex backgrounds. These point clouds are utilized to initialize two sets of 3D Gaussians (Frg-3DGS and Bkg-3DGS) aiming to represent foreground and background views. Both foreground and background views are then merged with a 2D learnable parameter map to reconstruct full views. In the second stage, we leverage the reconstruction ability developed in the first stage to impose the temporal constraints on the video diffusion model. This approach ensures the temporal consistency in the edited videos while maintaining high fidelity to the editing text prompt. We further propose a recursive and ensembled refinement by revisiting the denoising step and guidance scale used in video diffusion process with Video-3DGS. To demonstrate the efficacy of Video-3DGS on both stages, we conduct extensive experiments across two related tasks: Video Reconstruction and Video Editing. Video-3DGS trained with 3k iterations significantly improves video reconstruction quality (+3 PSNR, +7 PSNR increase) and training efficiency (×$\times 1.9$, ×$\times 4.5$ times faster) over NeRF-based and 3DGS-based state-of-art methods on DAVIS dataset, respectively. Moreover, it enhances video editing by ensuring temporal consistency across 58 dynamic monocular videos. Project website is available here.
@@ 1 Introduction
Therefore, we first design the framework of Video-3DGS to address the complexities of dynamic monocular video scenes. One of the key distinctions of the proposed method lies in its ability to excel in “wilder” monocular video scenarios (e.g., videos with larger object motion). To this end, we initially simplify the video sequences with two decomposition strategies. First, spatial decomposition, powered by an off-the-shelf open vocabulary video segmentor, mitigates background clutter 111background containing complex regions that clutter the foreground objects (e.g., trees in the background).. Second, temporal decomposition breaks down the entire video sequence into multiple shorter video clips with overlapping frames between neighboring clips. Given these decompositions, we can effectively extract 3D points of masked foreground moving objects in each video clip by introducing MC-COLMAP. Meanwhile, the cluttered background is modeled with spherical-shaped random 3D points surrounding the pre-extracted 3D points of foreground moving objects. For each video clip, Video-3DGS utilizes two sets of 3D Gaussians (Frg-3DGS and Bkg-3DGS) to represent foreground and background 3D points, respectively. We additionally implement multi-resolution hash encoding (Müller, 2021) with deformation networks (Yang et al., 2023) on both Frg-3DGS and Bkg-3DGS, which can boost performance and efficiency. Last but not least, to obtain the final rendered 2D outputs from both 3D Gaussians, we adopt a straightforward and effective 2D learnable parameter map to merge the two rendered images (rendered views from Frg-3DGS and Bkg-3DGS), resulting in faithful video frame representations.
@@ 2.2.1 MC-COLMAP
where $\mathcal{V}^{f}$ and $class$ denote the extracted segmentation masks of foreground moving objects and the user-guided text prompt (a required input for the segmentation network to specify the target object), respectively. For example, in the top left of Figure 4, we can consistently extract ‘Swan’ across frames in monocular video using spatial decomposition strategy.
Temporal Decomposition The segmentation masks $\mathcal{V}^{f}$ for foreground objects, which span the entire video sequence, present a processing challenge for COLMAP due to their intricate motion dynamics. To mitigate this complexity, stemming from diverse motion patterns throughout the video sequence, we partition the video sequence into multiple shorter video clips. This division ensures that objects exhibit reduced motion within each clip, thus facilitating more manageable processing for COLMAP. Formally, we address this by splitting $\mathcal{V}^{f}$ into multiple $M$ clips $\{\mathcal{V}^{f}_{j}\}_{j=1}^{M}$.
Rather than evenly dividing the video sequence into $M$ clips, each containing $k$ frames, we introduce a progressive scheme to address potential failure cases, such as when the foreground object remains static within a clip, making the $k$ frames inadequate for point cloud extraction or registration in SfM. Specifically, we begin with the first clip initialized with the first $k$ frames (i.e., $\mathcal{V}^{f}_{1}=\{\mathbf{x}_{i}\}^{k}_{i=1}$). If the COLMAP function $R$ fails to process the clip (i.e., $status_{1}\neq$ ‘Success’), we iteratively include one additional consecutive frame into the current clip until COLMAP returns a ‘Success’ $status$. Subsequently, the next clip commences with the last frame of previous clip, also initialized with $k$ frames. This process continues until the entire video sequence is processed. We term the resulting pipeline as MC-COLMAP, which applies COLMAP in a Masked and Clipped manner. Our MC-COLMAP system (denoted as $R_{MC}$) yields multiple sets of masked and clipped 3D point clouds, along with their corresponding views from $M$ clips as demonstrated in the top right of Figure 4 and following equation:
@@
We validate the effectiveness of MC-COLMAP in Table 1 and Table 3, demonstrating that MC-COLMAP not only improves the success rate over the original COLMAP, but also provides better-initialized 3D points for constructing 3D Gaussians.
@@ 2.3 Video-3DGS (2nd Stage): Plug-and-Play Refiner for Editing Videos
Single-phase Refiner We employ pre-trained 3D Gaussians to preserve the original structure and ensure temporal consistency in video editing. Edited video frames, denoted as
$\{I^{edited}_{i}\}^{N}_{i=1},$ (6)
can be generated by any zero-shot video editor. However, these edits may introduce style inconsistencies or cause original objects to be changed. To address this issue, we fix the positional parameters $(x,r,s)$ and deformation parameters ($\delta$) of both Frg-3DGS and Bkg-3DGS. At the same time, we adjust the color value ($SH$) and opacity ($\sigma$) parameters to match the edited style. This update is achieved by minimizing the reconstruction loss between the rendered images $\{I^{merged}_{i}\}^{N}_{i=1}$ and edited frames $\{I^{edited}_{i}\}^{N}_{i=1}$ for updating corresponding $SH$ and $\sigma$ as follows:
@@
Adjusting the color values ensures that corresponding regions in the rendered images remain consistent within each clip. Furthermore, by refining overlapping frames between adjacent clips, we achieve smoother style transitions and improved temporal consistency throughout the video. Given that we adopt Video-3DGS following a single-phase of video editing, Video-3DGS (2nd stage) can be appropriately referred to as a single-phase refiner.
Recursive and Ensembled Refiner In the single-phase refiner, we update $SH$ and $\sigma$ by minimizing the reconstruction loss between the rendered images and the edited frames. However, when the off-the-shelf video editor uses an excessive number of denoising steps ($N_{d}$) with editing prompts, some edited frames may lose essential structural details. To counteract this, we split $N_{d}$ into multiple phases (with $N_{r}=$ 2 for efficiency) and provide the detailed explanation for each phase.
First Phase: we first apply the signle-phase refinement Equation 7 on the initially edited frames (e.g., after $N_{d}$/2 denoising steps). This produces a set of intermediate refined frames.
Initialization for the Next Phase: we use the intermediate edited frames from above first phase as input for DDIM inversion (as in TokenFlow (Geyer et al., 2023) and RAVE (Kara et al., 2023) or an image encoder (as in Text2Video-Zero (Khachatryan et al., 2023)) to initialize the subsequent phase.
Second Phase: Then, we complete the remaining $N_{d}$/2 denoising steps for further video editing and follow with a final single-phase refinement on the newly edited frames.
We refer to this overall procedure as recursive refinement with multiple phases. Additionally, the guidance parameters in the diffusion process (e.g., image guidance in Text2Video-Zero, text guidance in TokenFlow, and ControlNet guidance in RAVE) play a crucial role in the quality of the edits. As shown in Figure 3, different guidance scales can lead to significant variations. To reduce sensitivity to these scale changes and combine the benefits of multiple scales, we perform an ensemble update of $SH$ and $\sigma$ using edited videos generated at three different scales during each recursive phase. The ensemble of edits from one phase is then incorporated into the next, allowing each phase to benefit from previous refinements. For simplicity, we denote the 2nd-stage Video-3DGS with Recursive and Ensembled refinement as Video-3DGS (RE), as illustrated in Figure 5.
@@ 3.2 Video Editing
Importance of 3D Gaussians for Video Editing Task The primary goal of Video-3DGS is to enhance the temporal consistency of video editing by leveraging 3D Gaussian representations with a two-stage approach. Specifically, in the 1st stage, our Video-3DGS reconstructs videos using 3D Gaussians to establish temporal correspondence between frames within each clip and link neighboring clips with overlapping frames. It motivates us to prove the video reconstruction capability of Video-3DGS, showcasing its ability to preserve the content and structure of original raw video without need of any 3D task exploration. Afterwards, we are able to exploit this reconstruction ability for video editing tasks in the 2nd stage. We provide explicit evidence of 3D Gaussians’ suitability for temporal correspondence as shown in this video link, where we visualize the correspondence for the first five groups of 3D Gaussians. As shown in the visualization, Video-3DGS ensures the correspondences of similar areas across frames using 3D Gaussians of the same group. We can enhance this further by incorporating the visualization of pixel trajectories tracked by the 3D Gaussians. This enhancement is demonstrated in the provided video link.
@@
Quantitative Results We assessed Video-3DGS’s editing capability on top of those three zero-shot video editors in Table 2. Video-3DGS significantly enhances temporal consistency (WarpSSIM score) across various editing scenarios spanning three different datasets. Consequently, this improvement generally yields superior final video editing results ($Q_{edit}$). It proves that Video-3DGS generally provides robust temporal editing guidance to existing video editors. We also observe that our proposed recursive and ensembled refinement can further improve both temporal consistency and video editing results on all of video editors and datasets.
@@
4 Discussion & Limitations
While the proposed Video-3DGS exhibits remarkable performance in both video reconstruction and editing tasks, certain challenges persist.
In video reconstruction, our Video-3DGS still relies on the success of MC-COLMAP even though it is more robust version compared to original COLMAP. Following are scenarios when MC-COLMAP can struggle or fail. Significant Occlusions of Foreground Objects: When a foreground object becomes fully or partially occluded in consecutive frames, MC-COLMAP’s decomposition process can struggle to track and re-identify the object. Excessive Motion Blur: Rapid camera or object movements can induce motion blur, making it difficult for feature detection algorithms to match points across frames accurately. If the blurred regions prevent consistent feature extraction, the reconstruction pipeline may fail to build a coherent model for certain frames or produce large alignment errors. We further acknowledge that MC-COLMAP for Video-3DGS requires a preprocessing step for foreground object extraction using an off-the-shelf panoptic segmenter. However, this is notably simpler compared to the five-stage preprocessing pipeline of shape-of-motion Wang et al. (2024a), which involves mask estimation, metric depth, monocular depth, camera estimation, and 2D tracking. Furthermore, we studied the influence of different variants and settings of the off-the-shelf segmenter on video reconstruction quality to ensure robustness as we discuss more in “Ablations on Segmentation Modules” of Section D.1.
In video editing, unsatisfactory results may arise if the underlying video editor fails or if significant changes to object shapes are required, which contradicts the inherent property of 3D Gaussians prioritizing the preservation of original structure. As a future direction for Video-3DGS, we aim to develop an architecture capable of manipulating 3D Gaussians in a more flexible manner to accommodate desired shape changes. Nevertheless, our work represents a pioneering effort in extending 3DGS to enhance the temporal consistency of initially edited videos from any zero-shot video editor. We envision that this will inspire further research in a zero-shot video editing community.
Unlike other conventional neural representation papers, our work does not explore the task of novel view synthesis. Instead, our method focuses on video reconstruction and editing by encoding and fitting videos with 3DGS. We anticipate that our Video-3DGS will also serve as a fundamental framework for 4D novel view synthesis in future research.
@@ 5 Conclusion
Acknowledgement This work was supported in part by the Institute of Information and Communications Technology Planning and Evaluation (IITP) grant funded by the Korea Government (MSIT) (Artificial Intelligence Innovation Hub) under Grant 2021-0-02068. This research was also partially supported by the KAIST Cross-Generation Collaborative Lab Project.
@@
  • Appendix A provides the related works.
@@
  • Appendix G discusses broader impact.
@@ Appendix A Related Work
In this part, we introduce the literature related to the core part of Video-3DGS (1st stage): neural rendering for scene representation (Section A.1), 3D Gaussian Spatting-based representation (Section A.2), dynamic scene representation (Section A.3), and spatial temporal decomposition in neural rendering (Section A.4). Regarding to Video-3DGS (2nd stage), we briefly review the literature on Text-to-Image editing (Section A.5), and Zero-shot Text-to-Video editing (Section A.6).
@@ A.1 Neural Rendering for Scene Representation
Building upon the advancements in Neural Radiance Fields (NeRF) (Kerbl et al., 2023) for scene representation, another innovative approach that has gained attention is the utilization of 3D Gaussian distributions to model scenes. This methodology, often referred to as “Gaussian Scene Representation”, leverages the mathematical properties of Gaussian distributions to efficiently capture the volumetric density and color information of 3D scenes. Similar to the trend in NeRF, there have been several advancements in performance and efficiency. For example, some of them apply the module used in NeRF to 3D Gaussian-based view synthesis.  (Lee et al., 2024) leverages a compact MLP to adjust the covariance of 3D Gaussians, which can effectively reconstruct sharp details from blurry images and Mip-splatting (Yu et al., 2023) address alias issue in 3DGS as similar to  (Barron et al., 2021). Besides the advancements in quality,  (Fan et al., 2023) enhances the efficiency and compactness of 3D Gaussian Splatting for neural rendering by pruning insignificant Gaussians.
@@ A.4 Spatial Temporal Decomposition
Our proposed Video-3DGS decomposes the dynamic video spatially and temporally to represent original video effectively. Similarly, ProgressiveNeRF (Meuleman et al., 2023) proposes method of decomposing video into spatial and temporal components through progressively optimized local radiance fields. It represents a significant advancement in the field of robust view synthesis and large-scale scene reconstruction. The primary distinction between our method and the ProgressiveNeRF approach lies in how we manage camera poses and frame integration across video clips. While ProgressiveNeRF dynamically updates camera poses and continues to append frames until the camera’s pose extends to the boundary of a high-resolution, uncontracted space, our method treats each clip independently. We utilize overlapping frames solely to ensure consistency of 3D Gaussians corresponding to overlapping frames across neighboring clips, with each clip maintaining its unique camera poses. However, all clips can still be rendered onto the same image plane via a differentiable rendering with each distinctive 3D Gaussian and camera pose. Additionally, our process repeats until MC-COLMAP signals a ‘Success’ status marking a clear difference in the criteria used to determine the sufficiency of frames for each clip. Furthermore, ProgressiveNeRF uses several dependencies (e.g., depth estimation and optical flow) and many losses for better temporal decomposition. However, we only need segmentation masks and simple reconstruction losses (L1 and SSIM) as in the original 3DGS paper. We also provide extensive visual comparison with ProgressiveNeRF in this link.
A.5 Text-to-Image Editing
A.6 Zero-shot Text-to-Video Editing